# Non-Smooth Solutions of the Navier-Stokes Equation

### v5

@@ V Part III: Analyticity
FoiManRos04 , eq. (8.1). Here we write $(-\Delta)(u_{1}+iu_{2})=(-\Delta)u_{1}+i(-\Delta)u_{2}$. Due to the periodic boundary conditions, the Stokes operator has been reduced to $-\nu\Delta$. This simple identity reduces estimates of the complex term to estimates of the corresponding real terms, as in FoiManRos04 . We do not require analyticity for the Stokes operator, established in FoiManRos04 , in any case. We use the same splitting of the other complex terms in (33) into real and imaginary terms with individual estimates for the real and imaginary terms.
The complexified bilinear form $B$ is written as
$B(u,v)_{\mathbb{C}}=B(u_{1},v_{1})+B(u_{2},v_{2})+i[B(u_{1},v_{2})+B(v_{2},u_{1})]\ .$ (34)
The proof of analyticity for NSRI moments of order one is based on the energy conservation of Lemma III.2, with an entropy principle hypothesis and non-negativity of the turbulent dissipation rates. The result is a UV finite norm in $\mathcal{H}$ for $\nu_{t,NL}$ and $\nu_{t,Temp}$, with $\mathcal{H}$ continuity uniform in $t$ in a limit as the cutoffs are removed.
  • Proposition V.1 The real domain bounds of Lemma III.2 extend to the complex domain.
Proof. The real domain analysis is extended to the complex domain. As in FoiManRos04 , the complex domain analysis is repeated for the real and imaginary parts and the two bounds added. The result is still finite in $\mathcal{H}$, uniformly in $t$, with a larger but still finite norm.
@@
The proof is completed by the Lemma III.2 bounds in $\mathcal{H}$ on the turbulent dissipation terms that are uniform in $t$ as the cutoffs are removed. $\square$
@@
  • Proposition V.2 The time derivative of the turbulent energy $\nu_{t}$ vacuum expectation value is analytic, with a domain of analyticity at least as large as the domain of analyticity for $\nu_{t}$.
@@
VI Part IV: Blowup
VI.1 Blowup in Finite Time
Initial conditions which are not SRI exist within the subspace of spherical harmonics spaces $\ell\geq 2$.
Blowup is considered in the space of the energy spherical harmonics expansion, so that there is no need to develop the parallel enstrophy theory. Blowup is considered within a space of finite spherical harmonics coefficients.
Making use of the assertions, the proof of blowup is valid.
The numerical program of Hou23 documents a rapid near total blowup in enstrophy, followed by a slower blowup in the energy, to complete the blowup process.
We examine only the energy spherical harmonics, so that blowup in the energy is sufficient to show existence of non-smooth solutions. Within the energy spherical harmonics, it is sufficient to restrict the statistics to the single 3D mode $\ell=2$, as blowup in any turbulent statistic shows existence of non-smooth solutions. Analyticity in time for this sequence and its time derivative was shown in Propositions V.1 and V.2.
The weak solution is chosen having a non-zero LH weak entropy maximizing solution with non-zero second moment $NSRI$. This weak solution lies in $\mathcal{H}$ and defines a LH weak solution.
This choice of initial conditions necessarily results in blowup in a finite time.
@@ VI.1 Blowup in Finite Time
  • Theorem VI.1.1 Initial conditions exist for which every entropy maximizing solution of the Navier-Stokes equation has only a finite interval of regularity.
@@
Blowup occurs within the $\ell=2$ spherical harmonic space with $\nu_{t,NL}>0$. Initial conditions are chosen with this blowup mode strictly positive. These initial conditions are smooth spatially since they lie within the span of a finite number of eigenvalues of the Stokes operator.
@@
The discontinuity (and resulting non-smoothness of the solution) occurs in the blowup mode, which is the $\ell=2$ spherical harmonic expansion coefficient for the non-linear turbulent dissipation rate $\nu_{t}$. $\square$
@@
  • Theorem VI.1.2 Initial conditions exist for which every entropy maximizing solution of the Navier-Stokes equation is not smooth globally in time.
Proof. We must show that the solution is not smooth at the transition point marking the end of the interval of regularity. By definition, the end of the regularity region $t_{j}$ has the property that $u(t_{j})$ is not in $\mathcal{V}$. This is equivalent to saying that $\langle\Delta u(t),u(t)\rangle$ is discontinuous at $t=t_{j}$, which demonstrates the lack of smoothness globally in time. $\square$
@@
The discontinuity is in the energy dissipation, as it is converted from turbulent to viscous. A finite unit of turbulent energy dissipation has transitioned through a large range of small length scales, at or below the Kolmogorov scale, where viscosity is the primary dissipative mechanism, to become totally thermal energy dissipation.
VII Acknowledgements and Disclaimer

### v6

@@ 5 Analyticity
[13], eq. (8.1). Here we write $(-\Delta)(u_{1}+iu_{2})=(-\Delta)u_{1}+i(-\Delta)u_{2}$. Due to the periodic boundary conditions, the Stokes operator has been reduced to $-\nu\Delta$. This simple identity reduces estimates of the complex term to estimates of the corresponding real terms, as in [13]. We do not require analyticity for the Stokes operator, established in [13], in any case. We use the same splitting of the other complex terms in (31) into real and imaginary terms with individual estimates for the real and imaginary terms.
The complexified bilinear form $B_{\mathbb{C}}$ is written as sum of its real and imiganary parts
The proof of analyticity for NSRI moments of order one is based on the energy conservation of Lemma 3.2, with an entropy principle hypothesis and non-negativity of the turbulent dissipation rates. The result is a UV finite norm in $\mathcal{H}$ for $\nu_{t,NL}$ and $\nu_{t,Temp}$, with $\mathcal{H}$ continuity uniform in $t$ in a limit as the cutoffs are removed.
  • Proposition 5.1 The real domain bounds of Lemma 3.2 extend to the complex domain.
Proof. The real domain analysis is extended to the complex domain. As in [13], the complex domain analysis is repeated for the real and imaginary parts and the two bounds added. The result is still finite in $\mathcal{H}$, uniformly in $t$, with a larger but still finite norm.
@@
The proof is completed by the Lemma 3.2 bounds in $\mathcal{H}$ on the turbulent dissipation terms that are uniform in $t$ as the cutoffs are removed. $\square$
@@
  • Proposition 5.2 The time derivative of the turbulent energy $\nu_{t}$ vacuum expectation value is analytic, with a domain of analyticity at least as large as the domain of analyticity for $\nu_{t}$.
@@
6 Blowup
  • Definition 6.1 Initial data $u_{0}$ is called non-turbulent (or smooth) if its expansion in spherical harmonics basis functions is proportional to a mean value in the $m$ variables and a mean value in the $\ell$ variables.
    The data is otherwise turbulent. The data is strictly turbulent if it is orthogonal to the smooth data at each fixed $m$ and also orthogonal to the smooth data as a function of $\ell$.
The three dimensional spherical harmonics space defined by $\ell=1$ is used to illustrate this definition. In this space, the mean is the lowest energy state in this subspace. It is SRI. The two states orthogonal to it are strictly turbulent. This two dimensional space will be used to construct blowup solutions. This space is NSRI.
Any finite dimensional subspace is not orthogonal to the $\ell$ mean and is thus turbulent.
Making use of the assertions, the proof of blowup is will be shown to be valid.
The numerical program of [21] documents a rapid near total blowup in enstrophy, followed by a slower blowup in the energy, to complete the blowup process.
The weak solution is chosen as LH weak entropy maximizing solution with strictly turbulent initial conditions. This weak solution lies in $\mathcal{H}$ and defines a LH weak solution.
This choice of initial conditions is conjectured to result in blowup in a finite time. Restricted to $\ell=1$, we prove blowup.
@@ 6 Blowup
  • Conjecture 6.1 Each entropy maximizing solution of the Navier-Stokes equation with strictly turbulent initial conditions has only a finite interval of regularity. It is not globally smooth in time.
  • Theorem 6.1 Each entropy maximizing solution of the Navier-Stokes equation with strictly turbulent initial conditions in the spherical harmonics subspace $\ell=1$ has only a finite interval of regularity. It is not globally smooth in time.
@@
Blowup occurs within the $\ell=1$ spherical harmonic space with $\nu_{t,NL}>0$. Initial conditions are chosen with this blowup mode strictly positive. These initial conditions are smooth since they lie within the span of a finite number of eigenvalues of the Stokes operator.
@@
The discontinuity (and resulting non-smoothness of the solution) occurs in the blowup mode, which is the $\ell=1$ spherical harmonic expansion coefficient for the non-linear turbulent dissipation rate $\nu_{t}$.
We next show that the solution is not smooth at the transition point marking the end of the interval of regularity. By definition, the end of the regularity region $t_{j}$ has the property that $u(t_{j})$ is not in $\mathcal{V}$. This is equivalent to saying that $\langle\Delta u(t),u(t)\rangle$ is discontinuous at $t=t_{j}$, which demonstrates the lack of smoothness globally in time. $\square$
@@
The discontinuity is in the energy dissipation, as it is converted from turbulent to viscous. A finite unit of turbulent energy dissipation has transitioned through a large range of small length scales, at or below the Kolmogorov scale, where viscosity is the primary dissipative mechanism, to become totally viscous energy dissipation.
7 The Mean
The mean $\mathcal{M}(u)$, defined in the vector spherical harmonic space $\mathcal{H}$ is a sum of three separate scalar spherical harmonics means. One is defined in the energy expansion and the other two are the $\pm$ oriented enstrophy expansions. The $\pm$ vorticity orientation reduces to two separate formulas for the mean $\mathcal{M}(u)$. The energy expansion has no dependance on vorticity, so that the energy mean can be determined independently of the spherical harmonics $m$ values.
7.1 Entropy Production Minimization
From the solid harmonics expansion of a weak solution $u$ of the Navier-Stokes equation or of any polynomial function of $u$, we note that the total entropy is equal to the sum of two individual entropies, defined by the $r$ and the $\ell$ solid harmonics expansion variables.
A Legendre transform relates $\ell$ energy dissipation to the $\ell$ entropy.
All proofs are of the form of uniform properties of cutoff quantities. The Legendre transforms are established in the cutoff case. Thus we can reason in terms of energy dissipation.
The Minimum Entropy Principle
The minimum entropy principle (when valid for a specific solution) states that a nonnegative polynomial function of the solution has a minimum rate of entropy production relative to the same polynomial function of any other weak solution of the Navier-Stokes equation with the same initial conditions.
A Key Distinction
The key distinction between the two (opposite) cases of entropy production maximizing vs. minimizing is the fact that for an entropy production minimizing solution, the turbulent dissipation rate $\nu_{t}$ is zero.
The blowup in Fig. 1 changes its character: the diagonal sloping line generated by the time derivative is now horizontal, and never intersects the horizontal line below it defined by the energy. Thus there is no end to the duration of the strong solution, and the mean solution is smooth for all space and time.
  • Remark 7.1.1 The minimum entropy principle is used in this paper to study Navier-Stokes solutions and also nonnegative polynomial functions of $\nu_{t,NL}$ and $\nu_{t,Temp}$.
    The spherical harmonics energy defined by $\ell$ is intrinsically nonnegative. A nonnegative entropy principle polynomial for $\ell$ must have nonnegative even coefficients and zero odd coefficients.
  • Theorem 7.1.1 Any element of $\mathcal{H}$ is expressed uniquely in terms of its spherical and solid harmonics expansion variables, $r,\ell,m$.
    The mean $\mathcal{M}(u)$ is the sum (normalized by division by the number of terms in the sum) of
    1. 1.
      the $r$ configuration space (magnitude equal ±$\pm 1$) mean,
    2. 2.
      the loop expansion phase space mean defined at fixed spatial coordinates,
    3. 3.
      the $\ell$ terms defined at fixed $t$,
    4. 4.
      the $m$ mean defined at fixed $t$ and fixed $\ell$.
    The mean has entropy zero.
The proof is placed in Sec. 9.
7.2 Regularity of the Mean
For a nonnegative polynomial function $p(u)$ of a weak solution of the Navier-Stokes equation, its average value $\mathcal{M}(|p(u)|)$ is defined as the ratio $\langle|p(u)|\rangle/\langle 1\rangle$. In simple terms it is the sum of values divided by the number of terms in the sum.
Here we study $\mathcal{M}(\nu_{t,Temp})$ and $\mathcal{M}(\nu_{t,NL})$. Products of distributions are not a concern, as we introduce cutoffs (so that all distributions become smooth functions which can be multiplied) and then prove results that are uniform as the cutoffs are removed.
The complete mean is the sum of the means defined by the spherical harmonic $r,\ell$ means.
  • Theorem 7.2.1
    $\mathcal{M}(\nu_{t,Temp})=0\in L^{4/3}(0,T;\mathcal{V}^{\prime})\subset\mathcal{S}^{\prime}$ (32)
    $\mathcal{M}(\nu_{t,NL})=0\in L^{4/3}(0,T;\mathcal{V}^{\prime})\subset\mathcal{S}^{\prime}$ (33)
  • Theorem 7.2.2
    $\nu_{t,Temp}(\mathcal{M}(\nu_{t,Temp}))=0\in\mathcal{S}^{\prime}$ (34)
    $\nu_{t,NL}(\mathcal{M}(\nu_{t,NL}))=0\in L^{4/3}(0,T;\mathcal{H}_{-3})\subset{S}^{\prime}\ .$ (35)
Intuitively, the evaluation is zero because the turbulent fluctuation of a turbulent fluctuation is zero.
Proofs are placed in Sec. 9.
8 The Mean is a Smooth Solution of the Navier-Stokes Equation
The main result of this section is the construction of smooth solutions of the Navier-Stokes equation with smooth non-turbulent initial data. The smooth solution is the mean of a weak solution. Its entropy is zero, so that it can also be characterized as an entropy production minimizing solution. We show that
  1. 1.
    Initial conditions: The mean of a weak solution of the Navier-Stokes equation with initial data $u^{0}\in\mathcal{H}(\mathbb{T}^{3})$ (the space of $L_{2}$ velocity fields defined on the cube $\mathbb{T}^{3}$ with periodic boundary conditions) the initial conditions are assumed in a continuous fashion, in the sense that
    $\mathcal{M}(u^{0})=\lim_{t\rightarrow 0}\mathcal{M}(u(t))\ .$ (36)
    in the topology of $\mathcal{H}(\mathbb{T}^{3})$.
    All vector fields are assumed to be divergence-free.
  2. 2.
    The mean of a weak solution of the Navier-Stokes equation is a strong solution. It is smooth for all time.
  • Theorem 8.1 Assume smooth initial conditions $u^{0}\in C^{\infty}$.
    The mean is smooth.
    The mean assumes its initial conditions (36).
    The mean of a weak solution $u$ of the Navier-Stokes equation in $\mathcal{H}$ is a solution of the Navier-Stokes equation in $\mathcal{H}$.
Proof. By Theorem 7.2.1, $\mathcal{M}(\nu_{t,NL}(u))=0\in\mathcal{H}_{-2}\subset\mathcal{S}^{\prime}$. and $\mathcal{M}(\nu_{t,Temp}(u))=0\in\mathcal{H}_{-2}\subset\mathcal{S}^{\prime}$.
To find the equation which the mean satisfies, we reason directly. $u$ satisfies the equation
$\partial_{t}u+\nu_{t,Temp}+\nu_{t,NL}(u)=\nu\Delta u\in\mathcal{S}^{\prime}\ .$ (37)
Since the mean $\mathcal{M}$ is linear, it is applied to each of the terms of (37) individually, in the space $\mathcal{V}^{\prime}\subset\mathcal{S}^{\prime}$. By Theorem 7.1.2 the mean is a solution of the Navier-Stokes equation in $\mathcal{V}^{\prime}$. $\mathcal{M}(\nu_{t,NL}(u))$ and $\mathcal{M}(\nu_{t,Temp}(u))$ are each the 0 element of $\mathcal{S}^{\prime}$.
By Theorem 7.2.1, there remains the equation
$\partial_{t}\mathcal{M}(u)=\nu\Delta\mathcal{M}(u)\ ,$ (38)
is valid in $\mathcal{V}^{\prime}\subset\mathcal{S}^{\prime}$.
This is the equation is also the heat equation. It is well known that the solutions in $\mathcal{S}^{\prime}$ of the heat equation with $\mathcal{S}$ initial conditions are smooth, belonging to $\mathcal{S}$, to prove the first assertion.
Solutions of the heat equation with initial data $u^{0}\in\mathcal{H}$ satisfy continuity of initial data (36). This statement is applied not to the weak solution $u$ but to $\mathcal{M}(u)$, with assumed unrestricted smooth initial data $\mathcal{M}(u^{0})$. Thus we assume $\mathcal{M}(u^{0})\in\mathcal{S}$, and the initial conditions (36) are assumed, the second assertion.
To prove the third assertion, recall the 0 evaluation in $\mathcal{S}^{\prime}$ of $\nu_{t,NL}(\nu_{t,NL}(\mathcal{M}{)u})$ and of $\nu_{t,Temp}(\mathcal{M}(\nu_{t,Temp}))$ according to Theorem 7.2.2.
Thus (38) is the Navier-Stokes equation in $\mathcal{S}^{\prime}$.
As inferred above, $\mathcal{M}(u)\in\mathcal{S}$ is smooth, so that (38) is the Navier-Stokes equation in $\mathcal{S}$. $\square$.
9 Appendix
9.1 Theorem 7.1.1
  • Theorem 9.1.1 The minimum entropy solution of the Navier-Stokes equation is unique. It is an element of $\mathcal{V}^{\prime}$.
Proof: By items 2,3, $\nu_{t,Temp}\in\mathcal{V}^{\prime}$ and $\nu_{t,NL}\in\mathcal{V}^{\prime}$.
By general theories of positive definite forms, uniqueness of a minimum is a result of strict convexity.
Strict convexity of the energy dissipation is a result of the viscous dissipation term $\nu\Delta u$, which sets a strictly positive lower bound on the dissipation as $\nu\Delta u$, with a gradient in $\mathcal{V}^{\prime}$. This completes the proof for a solution in $\mathcal{V}^{\prime}$. $\square$
Proof of Theorem 7.1.1 A mean is not a random variable. It is the expectation value of a random variable.
The first statement is elementary.
The mean of any of the summands 1 – 4 in the statement of the theorem is defined as
$\frac{\sum\mathrm{summand}}{\mathrm{number}\;\mathrm{of}\;\mathrm{terms}}\ .$ (39)
In the cutoff cases where this formula is used, the sum contains only a finite number of terms.
The $r$ phase space entropy production and $r$ phase space energy dissipation are evaluated at fixed spatial coordinates while the $\ell$ entropy production and energy dissipation are evaluated with the time coordinate fixed.
The $r$ configuration space dissipation is a sum of the energy dissipation for the two solid harmonic solutions, regular and irregular at the origin. The solution irregular at the origin is chosen as minimizing the energy dissipation. Uniqueness of this (one dimensional) minimum dissipation state follows.
The $r$ conjugate momentum space is the expansion space of the conjugate momenta $\pi$, called the loop expansion. Uniqueness of the entropy minimizing loop expansion variable is a consequence of Theorem 9.1.1.
The volume of the entropy minimizing solution is its $L_{1}$ expectation value, also equal to its $L_{1}$ norm. The volume of the solution divided by its $L_{1}$ norm is 1. With the unity temperature normalization, the entropy is the $L_{1}$ norm of this ratio.
By the definition of the entropy, the $r$ entropy of the mean is $\log 1=0$.
Next we specialize to a fixed time, so that the spherical harmonics provides a complete expansion of the mean of the $\ell,m$ variables.
The denominator in (39) is the number of energy dissipation terms in the numerator. We consider the spherical harmonic basis $Y_{\ell,m}$ and a partial mean $\mathcal{M}_{\ell}$ defined in terms of the sum over $\ell$.
All steps as in the $r$ case can be followed. The energy dissipation is intrinsicly nonnegative, to reduce the solution space to one dimension. To arrive at a one dimensional space, we use Theorem 7.1.2, which states that the $\ell$ entropy minimum is unique, or equivalently, that the $\ell$ minimum energy dissipation state is unique.
(Related analysis focused on entropy production maximization will be presented in joint work with H. Said and G.-Q. Chen.)
The conclusion is then the same as in the $r$ case. The entropy of the $\ell$ mean $\mathcal{M}_{\ell}(u)$ is zero.
The plus sign is chosen since the vacuum expectation value is nonnegative. The rest of the proof follows as in the $r$ case.
Thus
$\langle M(u)\rangle=\langle\mathcal{M}_{\ell}(u)\rangle_{\mathcal{H}}=1\ .$ (40)
Since the volume of $\langle\mathcal{M}_{\ell}(u)\rangle_{\mathcal{H}}=1$, the entropy (the log of this volume) of $\mathcal{M}_{\ell}=0$.
Entropy is always 0 or larger. Thus the mean is entropy minimizing. $\square$
  • Corollary 9.1.1 The mean satisfies the minimum entropy principle.
Proof. Since $\mathcal{S}$ and $\mathcal{S}^{\prime}$ are closed under multiplication, the polynomial powers needed here are automatic. Thus the proof follows as in Theorem 7.1.1. $\square$
9.2 Proof of Theorem 7.2.1
By items 2,3, $\nu_{t,Temp}\in\mathcal{V}^{\prime}$ and $\nu_{t,NL}\in\mathcal{V}^{\prime}$. The mean of $\nu_{t,Temp}$ is at least as regular as $\nu_{t,Temp}$. Likewise, the mean of $\nu_{t,NL}$ is at least as regular as $\nu_{t,NL}$.
To justify the value 0 for the mean, consider item 2 of Sec. 2.6 for (32) and item 3 for (33).
The time derivative $\partial_{t}u$ occurs exclusively in the $r$ expansion. We assert that the $r$ terms contributing to $\mathcal{M}(\nu_{t,Temp})$ sum to zero in $\mathcal{V}^{\prime}$.
The nonlinear term $-\langle u_{i}\partial_{t}u_{k}\rangle$ occurs only in the $r$ solid harmonics expansion. The term is evaluated at fixed spatial coordinates. As a result of Lemma 3.2, the energy in the $r$ loop expansion is conserved up to the amount dissipated by the viscous term $\Delta u$. This conservation is preserved by taking a mean. Thus the $r$ loop expansion terms contributing to $\mathcal{M}(\nu_{t,temp})$ sum to zero in $\mathcal{V}^{\prime}$.
The $r$ term (32) selects the solid spherical harmonics configuration space solution which is regular at the origin.
The proof of (32), based on the $r$ configuration space and the $r$ loop expansion is complete.
Next, we prove (33).
The nonlinear term $(u\cdot\nabla)u$ occurs exclusively in the $\ell,m$ spherical harmonics expansion. This expansion is applied to the solution considered at a fixed time. At fixed time, the time derivative and the $r$ expansion are removed from item 3 of Sec. 2.6. As a result of Lemma 3.2 the energy (in the $\ell,m$) expansion is conserved exactly up to the amount dissipated by the viscous term $\Delta u$. This conservation is preserved by taking a mean.
Assuming this, the $\ell,m$ terms contributing to $\mathcal{M}(\nu_{t,NL})$ sum to zero.
This term is nonnegative, which was to be demonstrated. The quadratic polynomial $\nu_{t,NL}$ has a nonnegative quadratic coefficient, to complete the proof of (33). $\square$
9.3 Proof of Theorem 7.2.2
To prove (34) analyticly, consider
$\begin{split}&\nu_{t,Temp}(\mathcal{M}(\nu_{t,Temp}))=\nu_{t,Temp}[\mathcal{M}\langle(-u_{i}\partial_{t}u)\rangle]\\ &=\langle\langle u_{j}\partial_{t}u\rangle\mathcal{M}(u_{i}\partial_{t}u)\rangle\\ &=\langle u_{j}\partial_{t}u\rangle\langle\mathcal{M}(u_{i}\partial_{t}u)\rangle=0\ .\\ \end{split}$ (41)
Next we prove (35) analytically. Consider
$\begin{split}&\nu_{t,NL}\mathcal{M}(\nu_{t,NL})=((u\cdot\nabla)\cdot\nabla u)\mathcal{M}((u\cdot\nabla)u)\\ &((u\cdot\nabla)\cdot\nabla u)((u\cdot\nabla)(u\cdot\nabla)u)\ .\\ \end{split}$ (42)
and focus on the RHS two rightmost $\nabla$. The first and second of the four $\nabla$ describe fluctuations in the rightmost third and fourth $\nabla$.
Direct evaluation of these two rightmost $\nabla$ of (LABEL:eq:mean-mean) identifies the result as the zero element in the space of of cubic polynomials in elements of the space $L^{4/3}(0,T;\mathcal{H}_{-2})$.
The left two $\nabla$, acting on this zero element, is the zero element in the space of quintic polynomials in the still more singular space $L^{4/3}(0,T;\mathcal{H}_{-4})$. Thus $\nu_{t,NL}(\nu_{t,NL})=0\in\mathcal{S}^{\prime}$. This completes the proof of (35). $\square$
10 Acknowledgements and Disclaimer